An express train and a a local train start out from the same point and travel in opposite directions. The express train travels twice as fast as the local train. If they are 480 miles apart, after 4 hours, what is the average speed of each train?
Answer by mducky2(62)

One equation might be particularly useful:
Speed * Time = Distance

We can shorten these into variables we need:
speed of the express train = se
speed of local train = sl

Using the information from the problem. Both trains traveled the same time. We also know the total distance:
time = 4 hours
total distance = 480 miles

Now let's use the information we know. Since the express train travels twice as fast as the local train, we can set the speed of the express train in terms of the local train:
se = 2l

Let's plug it all back into the original equation:
Speed * Time = Distance
(se + sl) * 4 = 480
(se + sl) = 120

Since se = 2l
2sl + sl = 120
3sl = 120
sl = 40 mph

The average speed of the local train is 40 mph.

Now we can solve for the speed of the express train.
se = 2l
se = 2(40)
se = 80 mph

The average speed of the express train is 80 mph.